OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively.

OpenStudy (anonymous):

@satellite73 Help please!

OpenStudy (anonymous):

your in good hands here

OpenStudy (anonymous):

Find the value of the variable. If your answer is not an integer, leave it in simplest radical form. 

OpenStudy (psymon):

That just makes me think of the statefarm commercial.

OpenStudy (anonymous):

that's ALL STATE

OpenStudy (psymon):

Right, right....my bad D:

OpenStudy (anonymous):

like a good neighbour, @SithsAndGiggles is there

OpenStudy (psymon):

:D

OpenStudy (anonymous):

haha right

OpenStudy (anonymous):

jeez i wonder what he is writing ...

OpenStudy (anonymous):

can either one of yall help me with this?

OpenStudy (anonymous):

i think it's been like 5+ minutes lol i hope it's good :D

OpenStudy (anonymous):

@ThatHaley_Girl post in another thread and i will answer it

OpenStudy (psymon):

You should post it as it's own question @ThatHaley_Girl

OpenStudy (psymon):

He could've fallen asleep at the keyboard and his face is just on the keys.

OpenStudy (anonymous):

lol yeah

OpenStudy (anonymous):

\[a, ar, ar^2, ar^3, ar^4\] \[a, -12, ar^2, ar^3, 768\] so \[\frac{ar^4}{ar}=r^3=\frac{768}{-12}=-64\] now find \(r\) then \(a\)

OpenStudy (anonymous):

This might be a roundabout method, but here goes. A geometric sequence has the form \(\left\{a_n\right\}_{n=1}^k=\left\{br^{n-1}\right\}_{n=1}^k\). \(a_2=-12~\Rightarrow~br=-12\\ a_5=768~\Rightarrow~br^4=768\) Solving the first equation for \(b\): \[b=-\frac{12}{r}\] Substituting into the second, \[\left(-\frac{12}{r}\right)r^4=768\\ -12r^3=768\\ r^3=64\\ r=4~~~\text{(excluding complex solutions)}\] Now you can solve for \(b\) by substituting this into either equation, then you get the general sequence \(a_n\).

OpenStudy (anonymous):

Sorry for the wait guys ;)

OpenStudy (anonymous):

yeah, but after all that \(r\) is negative right?

OpenStudy (anonymous):

\[r^3=-64\iff r=-4\]

OpenStudy (anonymous):

Ah, yes, I left out the minus sign! Thanks for pointing it out @satellite73

OpenStudy (anonymous):

hmmm my answer choices are these: an = 3 • (-4)n + 1 an = 3 • 4n - 1 an = 3 • (-4)n - 1 an = 3 • 4n so doesn't the first term have to be 3? @SithsAndGiggles

OpenStudy (anonymous):

we have \(r=-4\) and \(ar=-12\) telling you \(a=3\)

OpenStudy (anonymous):

Yes it is ^^^

OpenStudy (anonymous):

therefore it is \[3\times (-4)^{n-1}\]

OpenStudy (anonymous):

oh ok got it was confused for a sec there! Thanks a bunch guys

OpenStudy (anonymous):

as usual, it is C it is always C

OpenStudy (anonymous):

we only found \(r\) in all our rambling

OpenStudy (anonymous):

lol yep it's always C.

OpenStudy (psymon):

Just because of that, we need a mcq where its a. b. d. e. and no c option.